OpenStudy (anonymous):

If y=x^3+2x and dx/dt=5, find dy/dt when x=2. Give only the numerical answer. @freckles

OpenStudy (freckles):

can you find dy/dx ?

OpenStudy (anonymous):

yes y'=3x^2+2x @freckles

OpenStudy (freckles):

dy/dx=3x^2+2

OpenStudy (freckles):

\[\frac{dy}{dx}=\frac{dy}{dt} \cdot \frac{dt}{dx}\]

OpenStudy (freckles):

\[3x^2+2=\frac{dy}{dt} \cdot \frac{dt}{dx}\]

OpenStudy (freckles):

you are given dx/dt=5 so dt/dx=1/5

OpenStudy (freckles):

and this occurs when x=2

OpenStudy (freckles):

\[3(2)^2+2=\frac{dy}{dt} \cdot \frac{1}{5}\]

OpenStudy (freckles):

can you solve for dy/dt

OpenStudy (anonymous):

doesd dy/dt=70

OpenStudy (freckles):

if i did my basic algebra correctly that is what i got

OpenStudy (anonymous):

\[y=x^3+2x\] diff. with respect to t \[\frac{ dy }{ dt }=3x^2\frac{ dx }{ dt }+2\frac{ dx }{ dt }\] substitute the given values